### v3

# Freeze-in Gravitational Waves and Dark Matter in Warm Inflation

Abstract
Recent study [1] has suggested that warm inflation may be realized with Standard Model gauge interaction alone. Motivated by this framework, we investigate the gravitational wave spectrum and graviton-portal dark matter production through the freeze-in process generated during warm inflation scenarios. We perform a comparative analysis for different dissipation terms, focusing on their distinct gravitational wave signatures in the high-frequency regime. Our findings reveal qualitative and quantitative differences in the spectral behavior, offering a preliminary pathway for exploring inflationary and dark matter models through high-frequency gravitational wave signals.
1 Introduction
Cosmic inflation [2, 3, 4] can naturally explain the cosmic microwave background radiation and the large-scale structure observations, since the density perturbation of the inflation field can provide the primordial seeds of our Universe. Warm inflation (WI) [5] is one type of inflationary scenario that fundamentally alters the conventional picture of the early universe’s expansion by including dissipative effects and concurrent radiation production during inflation. Unlike the standard cold inflation (CI) model, where the universe supercools and reheating occurs from the late stage of inflation [6], WI maintains a thermal bath throughout the inflation process by dissipating the inflaton field’s energy into radiation. This continuous particle production modifies the dynamics of the inflation field and the generation of primordial density perturbations, offering possible solutions to challenges faced by cold inflation such as the “eta problem” and allowing for sub-Planckian field excursions. The conceptual foundation of WI integrates quantum field theory techniques to model dissipative processes, showing that interactions between the inflation field and other fields produce radiation simultaneously with inflationary expansion.
Although the concept of WI presents a compelling alternative to conventional cold inflationary scenarios, constructing a theoretically consistent WI model has proven challenging. Early formulations of WI relied on scalar couplings between the inflaton field and the particles comprising the thermal bath. However, subsequent analyses revealed that scalar interactions predominantly induce thermal corrections to the inflaton potential, rather than generating significant dissipative friction. These thermal effects tend to hinder inflationary dynamics rather than facilitate them, complicating efforts to build viable WI models based on scalar field interactions [7].
Recently, K.V. Berghaus, M. Drewes, and S. Zell in ref. [1],demonstrate, for the first time, that WI can be consistently realized through Standard Model gauge interactions, and the CMB predictions of this model have been further investigated in [8]. This finding provides an intriguing framework for dissipative dynamics during inflation and establishes a concrete connection between early-universe cosmology and well-motivated particle physics interactions. Motivated by this idea in ref. [1],we investigate the production of gravitational wave (GW) originating from the scattering of thermal plasma constituents and the graviton-portal dark matter (DM) during the inflationary epoch and estimate their relic density. The GW produced during reheating in the standard cold inflation paradigm has been investigated before [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. In WI, the universe naturally transitions from the inflationary epoch to the radiation-dominated era without requiring a separate reheating phase, and gravitons can be continuously produced from plasma [22, 23, 24, 25, 26, 27]. The graviton and DM production mechanism closely resembles the freeze-in DM scenarios [28, 29, 30]. In this work, we mainly investigate the GW spectrum in the high-frequency regime and the associated DM production mediated by gravitons, within the framework of WI. Our analysis encompasses both a simplified version of the recently proposed WI scenario based solely on Standard Model gauge interactions [1], as well as the typical WI model considered in ref. [27].
This paper is organized as follows. In section 2, we briefly review the mechanism of WI. In section 3, we analyze the freeze-in mechanism of gravitons and DM in WI. In section 4, we calculate the GW spectrum for different WI models. In section 5, we also estimate DM production solely by gravitational interaction from Standard Model plasma, and study the correlation between the DM mass and the peak amplitude of GW spectra. Finally, a concise conclusion is given in section 6.
2 Warm inflation
During the WI, the evolution of the background is governed by the following equations:
$\ddot{\phi}+(3H+\Upsilon)\dot{\phi}+V^{\prime}(\phi)=0$ (2.1)
$\dot{\rho}_{r}+4H\rho_{r}=\Upsilon\dot{\phi}^{2}$ (2.2)
$H^{2}=\frac{8\pi}{3M_{\rm pl}^{2}}\left(\frac{\dot{\phi}^{2}}{2}+V(\phi)+\rho_{r}\right)$ (2.3)
where $\phi$ is the inflaton field, $H$ the Hubble parameter, $\rho_{r}$ the radiation energy density, and $M_{\rm pl}\simeq 1.221\times 10^{19}\penalty 10000\ \mathrm{GeV}$ the Planck mass, respectively. $\Upsilon$ is the dissipation coefficient which decides the energy transfer between the inflaton and the thermal bath. The distinction between the warm and cold inflation is that in the former model, the dissipation term causes the radiation energy not to vanish. In this work, we assume the inflation potential
$V(\phi)=\frac{\lambda}{4}\phi^{4}$ (2.4)
The radiation energy density with temperature $T$ is given by
$\rho_{r}=\frac{\pi^{2}}{30}g_{*}T^{4}$ (2.5)
where $g_{*}$ is the degrees of freedom of the thermal bath and we set $g_{*}=106.75$.
For convenience, one defines the dimensionless dissipation ratio $Q\equiv\frac{\Upsilon}{3H}$, which gives a measure of the strength of the dissipation effects in comparison to the Hubble expansion. Then, we can classify WI based on the parameter $Q$ as follows: for $Q>1$, strong dissipative regime; for $Q<1$, weak dissipative regime.
The slow-roll parameters in WI are modified to be
$\epsilon_{V}\equiv\frac{M_{\rm pl}^{2}}{16\pi(1+Q)}\frac{V^{\prime}}{V}\quad,\quad\eta_{V}\equiv\frac{M_{\rm pl}^{2}}{8\pi(1+Q)}\frac{V^{\prime\prime}}{V}$ (2.6)
And during inflation, we have
$\epsilon_{V}\simeq\epsilon_{H}\equiv-\frac{\dot{H}}{H^{2}}$ (2.7)
In this paper, we use $\epsilon_{H}$ to define the end of the inflation ($\epsilon_{H}=1$) and the onset of radiation domination ($\epsilon_{H}=2$).
The dissipation coefficient $\Upsilon$ has the general form
$\Upsilon\approx\kappa\frac{T^{a}\phi^{b}}{f^{a+b-1}}$ (2.8)
where $\kappa$ is a constant. $a$, and $b$ could be some numerical powers. $f$ could be the symmetry-breaking scale. There are some models that provide the specific form of $\Upsilon$. For example, refs. [31, 32, 1] realize the WI by introducing axion coupling to a pure Yang-Mills gauge group and give $\Upsilon\propto T^{3}$. Ref. [33] proposes a scenario which is similar to the little Higgs model, where $\Upsilon\propto T$.
In this work, we discuss these two cases separately, $\Upsilon\propto T$, $T^{3}$, and compare their corresponding predictions. For each case, we consider $Q_{\rm ini}=0.01$ and $Q_{\rm ini}=1$, where $Q_{\rm ini}$ denotes the value of $Q$ at the beginning of inflation. Using the WI2easy code [34], we have verified that both values satisfy the CMB constraints from Planck data [35].
3 Freeze-in production in warm inflation: gravitons and dark matter
In WI, gravitons are continuously produced from the plasma via particle scattering processes during and after inflation. Since gravitons are produced out of equilibrium and subject to Planck-scale suppression, their production mechanism is conceptually very similar to the ultra-violet DM freeze-in production mechanism [30]. And in ref. [28], the authors found that freeze-in-produced DM relic density could be significantly enhanced in WI compared with CI. In this section, we argue that this enhancing mechanism applies equally to freeze-in production of gravitons.
In this work, we only consider gravitons with sufficiently high momenta to ensure that they always remain within the horizon. Their evolution is governed by the Boltzmann equation
$\frac{d\rho_{\mathrm{GW}}}{dt}+4H\rho_{\mathrm{GW}}=\gamma$ (3.1)
where $\rho_{\rm GW}$ is the GW energy density. $\gamma$ corresponds to the production rate of graviton energy density from the SM thermal bath and it is a function of temperature $T$. Using dimensional analysis, we obtain
$\mathcal{\gamma}(T)\approx\mathcal{C}_{1}\frac{T^{7}}{M_{\rm pl}^{2}}+\mathcal{C}_{2}\frac{T^{9}}{M_{\rm pl}^{4}}+$... (3.2)
where $\mathcal{C}_{1}$, $\mathcal{C}_{2}$ are coefficients. The first term corresponds to the production of a single graviton. Since $T\ll M_{\rm pl}$ during WI as we will show later, we consider only the leading-order contribution. For double-graviton production, see refs. [30, 26].
Using $dN_{e}/dt=H$, the left-handed side of eq. (3.1) can be re-parametrized as
$\frac{d\rho_{\mathrm{GW}}}{dt}+4H\rho_{\mathrm{GW}}=He^{-4N_{e}}\frac{d(\rho_{\mathrm{GW}}e^{4N_{e}})}{dN_{e}}$ (3.3)
Then the solution to eq. (3.1) reads
$\rho_{\rm GW}=e^{-4N_{e}}\int_{N_{e}^{i}}^{N_{e}}e^{4N_{e}^{\prime}}\frac{\mathcal{\gamma}}{H}\ dN_{e}^{\prime}$ (3.4)
where we have assumed that the energy density of GW is zero at the onset of inflation.
Now, the production rate of the comoving GW energy density, $I_{\mathrm{GW}}=d(e^{4N_{e}}\rho_{\mathrm{GW}})/dN_{e}$, reads
$I_{\mathrm{GW}}(N_{e})=e^{4N_{e}}\frac{\mathcal{\gamma}}{H}=\mathcal{C}_{1}e^{4N_{e}}\frac{T^{7}(N_{e})}{M_{\rm pl}^{2}H(N_{e})}$ (3.5)
From eq. (3.5), during the inflation stage, we can see that $I_{\mathrm{GW}}$ is an exponentially increasing function. This behavior arises because the inflaton energy dominates the universe, as well as the temperature $T$ and the Hubble parameter $H$ remain approximately constant during the inflation stage.
And when the universe transits to the radiation domination stage, $I_{\mathrm{GW}}$ will be proportional to $e^{-N_{e}}$ since $T\propto e^{-N_{e}}$ and $H\propto e^{-2N_{e}}$. Therefore, $I_{\mathrm{GW}}$ should sharply peaked at some e-fold, $N_{\mathrm{peak}}^{\mathrm{GW}}$, which satisfies
$\left.\left[4+7\frac{d\ln{T(N_{e})}}{dN}-\frac{d\ln{H(N_{e})}}{dN}\right]\right|_{N_{e}=N_{\mathrm{peak}}^{\mathrm{GW}}}=0$ (3.6)
This is similar to the case of DM freeze-in process, where the DM yield reads [28]
$Y_{\mathrm{DM}}(N_{e})=\frac{45}{2\pi^{2}g_{*,s}}\frac{e^{-3N_{e}}}{T^{3}(N_{e})}\int_{N_{e}^{i}}^{N_{e}}I_{\mathrm{DM}}(N_{e}^{\prime})\ dN_{e}^{\prime}$ (3.7)
And $I_{\mathrm{DM}}$ is the production rate of the comoving DM number density
$I_{\mathrm{DM}}(N_{e})=e^{3N_{e}}\frac{T^{2n+4}(N_{e})}{\Lambda^{2n}H(N_{e})}$ (3.8)
where $\Lambda$ is the cutoff of the effective field theory which describes the interaction between DM and the radiation. For similar reason, the DM production should be sharply peaked at $N_{\mathrm{peak}}^{\mathrm{DM}}$, and it satisfies
$\left.\left[3+(2n+4)\frac{d\ln{T(N_{e})}}{dN_{e}}-\frac{d\ln H(N_{e})}{dN_{e}}\right]\right|_{N_{e}=N_{\mathrm{peak}}^{DM}}=0$ (3.9)
Figure 1: The evolution of $I_{\mathrm{DM}}$ and $I_{\mathrm{GW}}$ in WI with dissipation term $\Upsilon\propto T^{3}$ and $\Upsilon\propto T$, respectively. All curves are normalized to their respective maximum values. The blue line corresponds to $Q_{\mathrm{ini}}=0.01$ and the red line corresponds to $Q_{\mathrm{ini}}=1$. The black dashed line indicates the end of inflation ($\epsilon_{H}=1$), while the red and blue dashed lines correspond to the onset of radiation domination ($\epsilon_{H}=2$), respectively. Note that $I_{\mathrm{DM}}$ and $I_{\mathrm{GW}}$ both sharply peaked at some e-folding numbers in the post-inflationary epoch, prior to the onset of radiation domination. It can be also found that the reheating process is much more efficient for $\Upsilon\propto T^{3}$ than for $\Upsilon\propto T$.
Figure 1 shows the evolution of $I_{\mathrm{DM}}$ and $I_{\mathrm{GW}}$ in WI for both $\Upsilon\propto T^{3}$ and $\Upsilon\propto T$, under both $Q_{\mathrm{ini}}=0.01$ and $Q_{\mathrm{ini}}=1$. We set $n=2$ since this is the case of purely gravitational DM as we will discuss later. All curves are normalized to their respective peak values. The blue and red solid lines represent the cases with $Q_{\mathrm{ini}}=0.01$ and $Q_{\mathrm{ini}}=1$, respectively. The black dashed line marks the end of inflation, while the blue and red dashed lines indicate the onset of radiation domination for each scenario. Notably, both $I_{\mathrm{DM}}$ and $I_{\mathrm{GW}}$ exhibit obvious peaks at specific e-folding numbers during the post-inflationary epoch, preceding the transition to radiation domination.
Now, the DM yield can be expressed approximately as [28]
$Y_{\rm DM}(N_{e})\simeq\frac{45}{2\pi^{2}g_{*s}}\frac{e^{3(N_{\mathrm{peak}}^{\rm DM}-N_{e})}}{\Lambda^{2n}T^{3}(N_{e})}\Delta N_{\mathrm{peak}}^{\rm DM}\times\frac{T^{2n+4}(N_{\mathrm{peak}}^{\rm DM})}{H(N_{\mathrm{peak}}^{\rm DM})}\quad(N_{e}>N_{\mathrm{peak}}^{\rm DM})$ (3.10)
with $\Delta N_{\mathrm{peak}}^{\rm DM}\gtrsim 1$ being the half-width of $I_{\rm DM}(N_{e})$.
In CI, DM can be produced during the radiation-dominant epoch which starts at some initial time $N_{e}^{\rm RH}$ with the reheating temperature $T_{\mathrm{RH}}$. The DM yield then reads
$Y_{\rm DM}^{\mathrm{RH}}\simeq\frac{1}{4\sqrt{\pi}}\left(\frac{45}{\pi^{2}g_{*}}\right)^{\frac{3}{2}}\frac{1}{2n-1}\frac{M_{\rm pl}T_{\mathrm{RH}}^{2n-1}}{\Lambda^{2n}}$ (3.11)
In WI, the reheating temperature is defined as the plasma temperature when the universe becomes radiation-dominated ($\epsilon_{H}=2$). For identical reheating temperatures, we can compare the DM relic abundance from WI and CI,
$\frac{Y_{\rm DM}(N_{e})}{Y_{\rm DM}^{\mathrm{RH}}}\simeq(2n-1)\frac{I_{\rm DM}(N_{\mathrm{peak}}^{\rm DM})}{I_{\rm DM}(N_{e}^{\mathrm{RH}})}\Delta N_{\mathrm{peak}}^{\rm DM}$ (3.12)
Since $I_{\rm DM}(N_{\mathrm{peak}}^{\rm DM})\gg I_{\rm DM}(N_{e}^{\rm RH})$ and $\Delta N_{\mathrm{peak}}^{\rm DM}\gtrsim 1$, the relic abundance is significantly enhanced in WI compared with CI. Same argument should apply to the freeze-in GW. Analogously to freeze-in DM, it is useful to define the dimensionless quantity $\rho_{\rm GW}/s^{\frac{4}{3}}$ with $s=\frac{2\pi^{2}g_{*s}}{45}T^{3}$ the entropy density [22]
$\frac{\rho_{\rm GW}}{s^{\frac{4}{3}}}\simeq\left(\frac{2\pi^{2}g_{*s}}{45}\right)^{-\frac{4}{3}}\mathcal{C}_{1}\frac{e^{4(N_{\mathrm{peak}}^{\rm GW}-N_{e})}}{M_{\rm pl}^{2}T^{4}(N_{e})}\Delta N_{\mathrm{peak}}^{\rm GW}\times\frac{T^{7}(N_{\mathrm{peak}}^{\rm GW})}{H(N_{\mathrm{peak}}^{\rm GW})}\quad(N_{e}>N_{\mathrm{peak}}^{\rm GW})$ (3.13)
where $\Delta N_{\mathrm{peak}}^{\rm GW}\gtrsim 1$ is the half-width of $I_{\rm GW}(N_{e})$.
For CI with the same reheating temperature $T_{\mathrm{RH}}$, by integrating eq. (3.4) from $N_{e}^{\mathrm{RH}}$ to infinity we can obtain
$\bigg[\frac{\rho_{\rm GW}}{s^{\frac{4}{3}}}\bigg]_{\mathrm{CI},T_{\rm RH}}\simeq\left(\frac{2\pi^{2}g_{*s}}{45}\right)^{-\frac{4}{3}}\mathcal{C}_{1}\frac{T_{\rm RH}^{3}}{M_{\rm pl}^{2}H_{\rm RH}}$ (3.14)
where $H_{\rm RH}$ is the Hubble parameter at $N_{e}^{\rm RH}$.
The GW production in WI is therefore enhanced by a factor of
$\bigg[\frac{\rho_{\rm GW}}{s^{\frac{4}{3}}}\bigg]\bigg/\bigg[\frac{\rho_{\rm GW}}{s^{\frac{4}{3}}}\bigg]_{\mathrm{CI},T_{\rm RH}}\simeq\frac{I_{\rm GW}(N_{\rm peak}^{\rm GW})}{I_{\rm GW}(N_{e}^{\rm RH})}\Delta N_{\mathrm{peak}}^{\rm GW}$ (3.15)
We conclude that GW are primarily generated around $N_{\rm peak}^{\rm GW}$, with their production enhanced in WI. Moreover, we expect a more pronounced enhancement for the WI model with $\Upsilon\propto T$ than for the $\Upsilon\propto T^{3}$ model, as the more efficient reheating process in the latter case leads to a significantly smaller value of the ratio $I_{\rm GW}(N^{\rm GW}_{\rm peak})\big/I_{\rm GW}(N^{\rm RH}_{e})$.
4 Gravitational waves spectrum
Now we turn to the analysis of the GW spectrum. The graviton production rate density $\gamma$ takes the form [22, 23]
$\gamma(T)=\frac{4T^{7}}{M^{2}_{\rm pl}}\int\frac{d^{3}\hat{k}}{(2\pi)^{3}}\hat{\eta}(\hat{k},T)$ (4.1)
withk^=k/T\ $\hat${k}=k/T. Here $k$ represents the momentum of graviton. The $\hat{\eta}$ function takes into account all the Standard Model contributions and has been calculated in refs. [22, 23] using the leading-log approximation. The calculation including hard thermal loop resummation has been done in ref. [24]. Here we only consider the results of leading-log approximation [22],
$\hat{\eta}(\hat{k},T)=\frac{\hat{k}}{e^{\hat{k}}-1}\sum_{i}d_{i}\hat{m}_{i}^{2}(T)\ln\left(\frac{5}{\hat{m}_{i}(T)}\right)$ (4.2)
with
$d_{1}$ $=1,\quad d_{2}=3,\quad d_{3}=8,$ (4.3)
$\hat{m}_{1}^{2}$ $=\frac{11}{6}g_{1}^{2},\quad\hat{m}_{2}^{2}=\frac{11}{6}g_{2}^{2},\quad\hat{m}_{3}^{2}=g_{3}^{2}$
where $g_{1}$, $g_{2}$, $g_{3}$ are gauge couplings of the Standard Model $U(1)_{Y}$, $SU(2)_{L}$ and $SU(3)_{C}$ gauge groups, respectively.
The GW energy density is given by eq. (3.4) as
$\rho_{\mathrm{GW}}(N_{e})=\frac{2}{\pi^{2}M_{\rm pl}^{2}}e^{-4N_{e}}\int_{N_{e}^{i}}^{N_{e}}dN_{e}^{\prime}\ e^{4N_{e}^{\prime}}\frac{T^{7}}{H}\int d\hat{k}\ \hat{k}^{2}\hat{\eta}(\hat{k},T)$ (4.4)
where $N_{e}$ is chosen in the radiation-dominated era and long after the peak of $I_{\mathrm{GW}}$.
We can then calculate the GW energy spectrum
$\Omega_{\mathrm{GW}}=\frac{1}{\rho_{c}}\frac{d\rho_{\mathrm{GW}}}{d\ln k}$ (4.5)
where $\rho_{c}$ is the critical density of universe. We redshift the GW spectrum to the present time using
$\rho_{\gamma}=\rho_{\gamma,0}\left(\frac{a_{0}}{a}\right)^{4}\left(\frac{g_{*s,0}}{g_{*s}(N_{e})}\right)^{\frac{4}{3}}\quad,\quad\frac{k(N_{e})}{T(N_{e})}=\frac{k_{0}}{T_{0}}\left(\frac{g_{*s}(N_{e})}{g_{*s,0}}\right)^{\frac{1}{3}}$ (4.6)
where $g_{*s,0}=3.931$ is the present degrees of freedom of the entropy density. $\rho_{\gamma,0}$ is the current photon energy density.
Figure 2: The freeze-in GW spectrum of WI in the high frequency regime for different $Q_{\rm ini}$ and different dissipation terms. The dashed lines correspond to the freeze-in GW spectrum in CI produced during the radiation-dominated epoch with the same reheating temperature. The quartic coupling $\lambda$ is chosen to satisfy the CMB constraints. The upper panel and lower panel shows the GW spectrum for $Q_{\rm ini}=0.01$ and $Q_{\rm ini}=1$, respectively. Note the nearly identical spectral shapes and approximately equal peak frequencies, along with an enhanced peak amplitude compared to the CI case.
The GW spectrum today is given by
$\begin{split}\Omega_{\mathrm{GW},0}(k_{0})h^{2}=&\frac{2}{\pi^{2}M_{\rm pl}^{2}}\frac{\Omega_{\gamma,0}h^{2}}{\rho_{c}(N_{e})}\frac{g_{*}(N_{e})}{g_{\gamma}}\left(\frac{g_{*s,0}}{g_{*s}(N_{e})}\right)^{\frac{1}{3}}\times\\ &e^{-4N_{e}}\int_{N_{e}^{i}}^{N_{e}}dN_{e}^{\prime}\ e^{4N_{e}^{\prime}}\frac{T^{7}}{H}\left(\alpha(N_{e}^{\prime})\frac{k_{0}}{T_{0}}\right)^{3}\hat{\eta}\left[\alpha(N_{e}^{\prime})\frac{k_{0}}{T_{0}}\left(\frac{g_{*s}}{g_{*s,0}}\right)^{\frac{1}{3}},T\right]\end{split}$ (4.7)
where $\Omega_{\gamma,0}=5.38\times 10^{-5}$ , $g_{\gamma}=2$ and the dimensionless Hubble parameter $h=0.674$ [36]. We set $g_{*s}(N_{e})=g_{*}(N_{e})=106.75$. And we have
$\alpha(N_{e}^{\prime})=\frac{T(N_{e})e^{N_{e}}}{T(N_{e}^{\prime})e^{N_{e}^{\prime}}}$ (4.8)
Note that this function stays constant during the radiation-dominated era since $T(N_{e})\propto e^{-N_{e}}$. So in CI, as the GW are produced during the radiation-dominated era, the GW spectrum can be parametrized in the form of
$\Omega_{\mathrm{GW},0}^{\mathrm{CI}}(f)=\mathcal{A}\frac{(f/f_{\rm peak})^{4}}{e^{f/f_{\rm peak}}-1}$ (4.9)
where $f=\frac{k_{0}}{2\pi}$ is the present frequency. And the peak frequency can be derived from the maximum point of $\hat{k}^{4}/(e^{\hat{k}}-1)$,
$f_{\rm peak}^{\rm CI}\simeq 3.92\frac{T_{0}}{2\pi}\left(\frac{g_{*s,0}}{g_{*s}}\right)^{\frac{1}{3}}\approx 7.41\times 10^{10}\penalty 10000\ \mathrm{Hz}$ (4.10)
In WI, most GWs are generated around $N_{\mathrm{peak}}^{\mathrm{GW}}$, occurring after inflation and before the radiation-dominated era, as we have discussed. During this stage, $\alpha(N_{e}^{\prime})$ will slowly evolve over time, so the shape of GW spectrum will differ from $\Omega_{\mathrm{GW},0}^{\rm CI}$, and the peak frequency shifts. However, as we will show in our result, this change is very slight and $f_{\rm peak}^{\rm WI}\simeq f_{\rm peak}^{\rm CI}$, where $f_{\rm peak}^{\rm WI}$ is the peak frequency of freeze-in GW in WI.
In figure 2, we show the GW spectrum for different values of $Q_{\mathrm{ini}}$, both in the cases of $\Upsilon\propto T$ and $\Upsilon\propto T^{3}$. Additionally, the spectrum of freeze-in GW produced during the radiation-dominated epoch in CI are represented by dashed lines. The shapes of the spectra and the peak frequencies are nearly identical, while the peak amplitude is enhanced in WI. Note also that, as expected, this enhancement is much more pronounced in the $\Upsilon\propto T$ model than in the $\Upsilon\propto T^{3}$ case. We next delve into a comparison of the two models. When $Q_{\rm ini}=0.01$, the predictions of the two dissipative models are nearly indistinguishable. However, for $Q_{\rm ini}=1$, differences arising from the temperature dependence of the dissipation coefficients lead to a noticeable divergence: the peak amplitude is lower in the $\Upsilon\propto T$ model than in the $\Upsilon\propto T^{3}$ model. We also plot the sensitivity curves of LISA [37], TianQin [38], Taiji [39], BBO [40, 41], Cosmic Explorer (CE) [42], Einstein telescope (ET) [43, 44], and Resonant Cavities [45] for all panels.
5 Freeze-in production of purely gravitational dark matter
Purely gravitational DM [46, 47, 48] is one of the simplest possibilities for DM, since gravitational interaction is universal and model-independent (at least for sub-Planckian scales). It can be produced by the thermal bath mediated by gravitons, known as the gravitational DM freeze-in and it is a concrete realization of the DM freeze-in mechanism. The production of purely gravitational DM in early universe was carefully studied in CI [49, 50, 51, 52, 53], and was first considered in WI in ref. [54]. In this section, we focus on the correlation between the DM mass and the peak amplitude of the GW spectra as functions of the quartic coupling of inflaton field, applying our results from the last section.
We start with the Boltzmann equation of purely gravitational DM $\chi$ .
$\frac{dn_{\chi}}{dt}+3Hn_{\chi}=\langle\sigma v\rangle_{\chi}\big(n_{\chi}^{\rm eq}\big)^{2}$ (5.1)
where $n_{\chi}$ denotes the number density of Fermion DM $\chi$. And $n_{\chi}^{\mathrm{eq}}=\frac{3}{2}\frac{\zeta_{3}}{\pi^{2}}T^{3}$ is the DM number density in thermal equilibrium, where $\zeta_{3}=1.2\$is the value of Riemann-Zeta function.
The thermal averaged cross sections of all gravitational annihilation channels have been calculated in ref. [47]
$\langle\sigma v\rangle_{0\rightarrow\chi}=\frac{\pi m_{\chi}T}{2M_{\rm pl}^{4}}\left[\frac{4}{5}\frac{T}{m_{\chi}}+\frac{1}{5}\frac{m_{\chi}}{T}-\frac{1}{5}\frac{m_{\chi}}{T}\frac{K_{1}^{2}}{K_{2}^{2}}+\frac{2}{5}\frac{K_{1}}{K_{2}}\right]$ (5.2)
$\langle\sigma v\rangle_{\frac{1}{2}\rightarrow\chi}=\langle\sigma v\rangle_{1\rightarrow\chi}=\frac{4\pi m_{\chi}T}{M_{pl}^{4}}\left[\frac{6}{5}\frac{T}{m_{\chi}}+\frac{2}{15}\frac{m_{\chi}}{T}-\frac{2}{15}\frac{m_{\chi}}{T}\frac{K_{1}^{2}}{K_{2}^{2}}+\frac{3}{5}\frac{K_{1}}{K_{2}}\right]$ (5.3)
[…]
$\langle\sigma v\rangle_{0\rightarrow\chi}=\frac{2\pi T^{2}}{5M_{\rm pl}^{4}}\quad,\quad\langle\sigma v\rangle_{\frac{1}{2}\rightarrow\chi}=\langle\sigma v\rangle_{1\rightarrow\chi}=\frac{24\pi T^{2}}{5M_{\rm pl}^{4}}$ (5.4)
Thus the total cross section gives
$\langle\sigma v\rangle_{\chi}=4\langle\sigma v\rangle_{0\rightarrow\chi}+45\langle\sigma v\rangle_{\frac{1}{2}\rightarrow\chi}+12\langle\sigma v\rangle_{1\rightarrow\chi}=\frac{1376\pi T^{2}}{5M_{\rm pl}^{4}}$ (5.5)
Now eq. (5.1) becomes
$He^{-3N_{e}}\frac{dn_{\chi}e^{3N_{e}}}{dN_{e}}=\beta\frac{T^{8}}{M_{\rm pl}^{4}}$ (5.6)
where $\beta=3096\zeta_{3}^{2}/5\pi^{3}$. Then the solution of DM number density can be expressed as
$n_{\chi}(N_{e})=e^{-3N_{e}}\int_{N_{e}^{i}}^{N_{e}}I_{\chi}(N_{e}^{\prime})dN_{e}^{\prime}$ (5.7)
where
$I_{\chi}(N_{e})=e^{3N_{e}}\beta\frac{T^{8}(N_{e})}{M_{\rm pl}^{4}H(N_{e})}$ (5.8)
and the DM yield $Y_{\chi}\equiv n_{\chi}/s$ reads
$Y_{\chi}(N_{e})=\frac{45}{2\pi^{2}g_{*s}}\frac{e^{-3N_{e}}}{T^{3}(N_{e})}\int_{N_{e}^{i}}^{N_{e}}I_{\chi}(N_{e}^{\prime})dN_{e}^{\prime}$ (5.9)
We show the evolution of $Y_{\chi}$ in figure 3. The evolution of the Hubble parameter $H$, the plasma temperature $T$ and the parameter $Q$ are also given in figure 3. During the early stage of WI, although the inflaton energy dominates the universe, the thermal bath is still sustained with a nearly constant temperature and DM gradually accumulates. After the inflation stage, the universe transits to a radiation-dominated phase, where $T\propto e^{-N_{e}}$ and $H\propto T^{2}$.
Figure 3: The evolution of the Hubble parameter $H$, the plasma temperature $T$, the parameter $Q$, the radiation density parameter $\Omega_{r}$ and the DM yield $Y_{\chi}$.
Assuming that DM comprises only $\chi$ particles, their mass can be derived from
$\Omega_{\mathrm{DM},0}h^{2}=\frac{m_{\chi}Y_{\chi}(N_{e})s_{0}}{\rho_{c,0}}h^{2}\equiv 0.12$ (5.10)
where $s_{0}\simeq 2891.2\penalty 10000\ \mathrm{cm}^{-3}\$and $\rho_{c,0}\simeq 4.78\times 10^{-6}\penalty 10000\ \mathrm{GeV}\cdot\mathrm{cm}^{-3}$ [36] are the entropy density and the critical energy density today, respectively.
Figure 4: The correlation between DM mass and the peak amplitude of GW spectra for different $Q_{\rm ini}$ and dissipation terms. The colored bar represents the peak amplitude of GW spectra fixing the observed DM relic density. In the upper panel where $\Upsilon\propto T$, we found that in the case of $Q_{\mathrm{ini}}=1$, the DM mass does not satisfy the condition $m_{\chi}\ll T$ for $\lambda<2.6\times 10^{-16}$.
In figure 4, to satisfy the observed DM relic density, we show the correlation between DM mass and the peak amplitude of GW spectra for different $Q_{\rm ini}$ and dissipation terms. For each point on the curves, we have checked that the condition $m_{\chi}\ll T$ is always satisfied at $N_{e}=N_{\rm peak}^{\rm DM}$, when most DM is produced. Compared with the case $\Upsilon\propto T^{3}$, it can be seen that varying $Q_{\rm ini}$ has a stronger impact on both the predicted DM mass and the peak amplitude of the GW spectra in the case of $\Upsilon\propto T$. Moreover, the DM mass is significantly more sensitive to the coupling $\lambda$ than the GW peak amplitude is. And as the DM mass increases, the peak amplitude of the GW spectra decreases in all cases studied.
For a given $Q_{\rm ini}$, the value of $\lambda$ can be fixed by matching the observed scalar spectrum amplitude $A_{s}\approx 2.1\times 10^{-9}$ [35, 34]. Once $\lambda$ is determined, we then obtain the corresponding DM mass as well as the peak amplitude of the GW spectrum. Table 1 and table 2 list the values of $\lambda$, the scalar spectral index $n_{s}$, the tensor-to-scalar ratio $r$, DM mass, peak frequency, and the peak amplitude of the GW signal for WI model with $\Upsilon\propto T$ and $\Upsilon\propto T^{3}$, respectively. Notably, the GW spectra show nearly identical peak frequencies with amplitudes of order $10^{-13}-10^{-12}$, consistent with our previous results. The DM mass, however, varies significantly across models, ranging from $10^{6}$ to $10^{10}$ GeV.
$Q_{\mathrm{ini}}$ $\lambda$ $n_{s}$ $r$ $f_{\mathrm{peak}}(\mathrm{Hz})$ $\Omega_{\mathrm{peak}}h^{2}$ $m_{\chi}(\mathrm{GeV})$
0.01 $5.26\times 10^{-15}$ $0.964$ $8.05\times 10^{-3}$ $6.31\times 10^{10}$ $5.00\times 10^{-12}$ $9.44\times 10^{6}$
1 $3.55\times 10^{-15}$ $0.969$ $3.75\times 10^{-4}$ $6.03\times 10^{10}$ $3.34\times 10^{-13}$ $1.38\times 10^{10}$
Table 1: Model predictions for WI with $\Upsilon\propto T$: scalar spectral index $n_{s}$, tensor-to-scalar ratio $r$, peak frequency $f_{\mathrm{peak}}$, peak amplitude $\Omega_{\mathrm{peak}}h^{2}$, and DM mass $m_{\chi}$
$Q_{\mathrm{ini}}$ $\lambda$ $n_{s}$ $r$ $f_{\mathrm{peak}}(\mathrm{Hz})$ $\Omega_{\mathrm{peak}}h^{2}$ $m_{\chi}(\mathrm{GeV})$
0.01 $5.52\times 10^{-15}$ $0.963$ $8.87\times 10^{-3}$ $6.61\times 10^{10}$ $6.99\times 10^{-12}$ $5.88\times 10^{6}$
[…]
Table 2: Model predictions for WI with $\Upsilon\propto T^{3}$: scalar spectral index $n_{s}$, tensor-to-scalar ratio $r$, peak frequency $f_{\mathrm{peak}}$, peak amplitude $\Omega_{\mathrm{peak}}h^{2}$, and DM mass $m_{\chi}$
6 Conclusions
Motivated by the recent WI model with solely Standard Model gauge interactions [1], we have investigated the GW spectrum in the high frequency regime generated via the freeze-in mechanism in two types of WI models. We find that in WI, the freeze-in production of GW shows similar features to the DM freeze-in production, with both being enhanced compared with CI. Moreover, the shapes of GW spectra and the peak frequencies ($\sim 10^{10}\penalty 10000\ \rm Hz$) are nearly identical across different WI models, while the peak amplitude of the GW spectra today can be of the order $10^{-13}-10^{-12}$. We also studied the graviton-portal DM freeze-in scenario within two classes of WI framework. Under the constraint of reproducing the observed DM relic abundance, we have identified a correlation between the DM mass and the quartic coupling of the inflation field, and determined the corresponding peak amplitude of the GW signal. Notably, the predicted DM mass exhibits significant variation across different models. These findings provide a theoretical basis for probing such WI scenarios and DM production mechanisms through future GW observations.
Acknowledgments
This work is supported by the National Natural Science Foundation of China (NNSFC) under Grant No.12475111. and No.12205387.
Note added
While this work motivated by recent WI with solely Standard Model gauge interactions [1] was nearing completion, ref. [27] appeared on arXiv, which studied the thermal gravitons from warm inflation.